# High-Performance Multi-Qubit System with Double-Transmon Couplers towards Scalable Superconducting Quantum Computers

### v1

@@ III.2.1 Two-qubit gate: CZ gate
To implement $\hat{U}_{CZ,12(23)}$ in the three-qubit system, we fix $\Theta_{\rm ex,R(L)}$ to $\Theta_{\rm Id,R(L)}$ and apply a optimized flux pulse $\Theta_{\rm ex,L(R)}(t)$ (see Appendix G.1). We can successfully implement 30-ns CZ gates with average fidelities of 0.9999 for both the qubit pairs, as summarized in Table  3. Comparing these fidelities with the corresponding ones in the isolated two-qubit subsystems, we find that the degradation of the CZ gate fidelity due to the increasing number of qubits from two to three is smaller than 0.0001 in the DTC architecture.
@@ III.2.2 Single-qubit gate: $\pi/2$ pulse
Next, we consider the single-qubit gate implementation as well. Our target is a $10$-ns $\pi/2$ pulse for each qubit. It leads to the following $\hat{U}_{\rm id}$:
$\hat{U}_{\pi/2,1}$ $=\hat{U}_{\pi/2}\otimes\hat{I}\otimes\hat{I},$ (16)
$\hat{U}_{\pi/2,2}$ $=\hat{I}\otimes\hat{U}_{\pi/2}\otimes\hat{I},$ (17)
$\hat{U}_{\pi/2,3}$ $=\hat{I}\otimes\hat{I}\otimes\hat{U}_{\pi/2},$ (18)
@@
$\hat{U}_{\pi/2}=\frac{1}{\sqrt{2}}\begin{pmatrix}1&-1\\ 1&1\\ \end{pmatrix}.$ (21)
The average gate fidelity of $\hat{U}_{\pi/2,i}$ is denoted by $\bar{F}_{\pi/2,i}$ for $i\in\{1,2,3\}$.
To implement $\hat{U}_{\pi/2,i}$, we fix $\Theta_{\rm ex,\mu}$ to $\Theta_{\rm Id,\mu}$ and apply a driving voltage to ${\rm Q}_{i}$. This operation is described by the following Hamiltonian (see Appendix A for the derivation):
$\hat{H}_{V}$ $=4\hbar\hat{\mathbf{n}}^{\rm T}W\hat{\mathbf{n}}+\alpha_{i}\hbar\sum_{j=1}^{7}W_{ij}\hat{n}_{j}+\hat{U}(\Theta_{\rm Id,L},\Theta_{\rm Id,R}),$ (22)
where $\alpha_{i}$ is a dimensionless parameter controlled by $V_{i}$ (see Appendix A). Applying optimized $\alpha_{i}(t)$ (see Appendix G.2), we can implement 10-ns $\pi/2$ pulses with average gate fidelities over 0.9999 as summarized in Table  3.
In this section, we have shown that DTCs can not only reduce undesired residual couplings to several kHz but also enable implementations of fast CZ gates and $\pi/2$ pulses with average fidelities of 0.9999 or higher, in the three-qubit system as well as in isolated two-qubit systems. These results indicate that the DTC architecture is scalable for highly detuned fixed-frequency qubits.
@@ IV.2 Hamiltonian
Note that almost all previous theoretical analyses of the STC reported in the literature have been based on the effective model yan2018tunable ; PRXSTC ; RiggettiSTC ; STCChinese ; ETH ; IBMSTC ; YanSTCCZ ; TheoriticalSTCcrosstalk2022 . Unlike them, we use the circuit model for higher accuracy, as well as the DTCs. To the best of our knowledge, this is the first estimation of $ZZ$ and $ZZZ$ couplings and also the first gate simulation based on the circuit model for systems with the STCs.
Figure 5: Energy levels and $ZZ$ coupling strengths of the systems in Fig. 4. (a) and (b) show eigenfrequencies $\omega_{Q_{1},Q_{2},Q_{3},C_{4},C_{5}}$ as functions of $\Theta_{\rm ex,L}$ and $\Theta_{\rm ex,R}$, respectively. The qubit frequencies of the three-qubit system, $\omega_{Q_{1},Q_{2},Q_{3}}$ $(Q_{i}\in\{0,1\})$, are highlighted by being colored and bold, and the ones of the two-qubit system, $\omega_{Q_{i},Q_{j}}^{\mu}$ ($Q_{i},Q_{j}\in\{0,1\}$ and $\mu\in\{{\rm L,R}\})$ (see Appendix D), are shown by scatter plots. Eigenfrequencies $\omega_{Q_{1},Q_{2},Q_{3},C_{4},C_{5}}$ other than the above are represented by thin curves. (c) and (d) represent $ZZ$ and $ZZZ$ couplings as functions of $\Theta_{\rm ex,L}$ and $\Theta_{\rm ex,R}$, respectively. Solid curves represent the ones of the three-qubit system and scatter plot represent $\zeta_{12(23)}$ in the system L(R). The Orange (blue) dotted horizontal line is $\zeta_{\rm Id,23(12)}$ in the system R(L). Note that $\Theta_{\rm ex,R}$ in the left column and $\Theta_{\rm ex,L}$ in the right column are fixed to $\Theta_{\rm Id,R}$ and $\Theta_{\rm Id,L}$, respectively.
@@ IV.3 Comparing conditions
The preferable parameter setting for the comparison between the STC and DTC architectures is non-trivial. This is because the STC and the DTC obey different operating principles, resulting in different parameter values desired for high performance. Here, in order to make the comparison as fair as possible, we set the parameters as shown in Table 4 , considering the following points. (1) We set $\omega_{i}$ and $C_{ii}$ for $i\in\{1,2,3\}$ to be the same as in Table 1, which means that qubit frequencies and anharmonicities are almost the same as in the DTC case. (2) Under this condition, we set $C_{ij}$ that exists in the two-qubit subsystems and $I_{c4{\rm B}(c5{\rm B})}/I_{c4\rm{A}(c5{\rm A})}$ to the almost same values as the corresponding ones in Ref. ETH, , and we also set $\omega_{4(5)}$ to the almost same value as the coupler idling frequency in Ref. ETH, . Reference ETH, experimentally implemented a high-performance CZ gate (with a gate time of 38 ns and a fidelity of 97.9%), using a parameter setting close to the above condition (1) , where qubits frequencies are 5.038 and 5.400 GHz and both qubits capacitances are 77.8 fF. Moreover, they also showed that the magnitude of the residual $ZZ$ coupling can be suppressed to 60–80 kHz by an STC, which is known to be one of the most $ZZ$-coupling suppressed results for highly detuned qubits where a fast CZ gate can be implemented. We will show that our numerical simulation reproduces their results well. (3) As for the rest $C_{ij}$ that does not exist in the two-qubit subsystems (e.g. $C_{13},C_{15}$, and $C_{45}$), we choose feasible values estimated by electromagnetic simulator ANSYS Q3DAnsys assuming the scale-up design of the circuit in Ref. ETH, .
@@ IV.4 Numerical results
IV.4.1 Idling point, $ZZ$ couplings, and $ZZZ$ coupling
Figure 5(a)[5(b)] and Fig. 5(c)[5(d)] show $\omega_{Q_{1},Q_{2},Q_{3},C_{4},C_{5}}$ and $\zeta^{\prime}s$ as functions of $\Theta_{\rm L(R)}$, respectively, when $\Theta_{\rm ex,R(L)}$ is fixed to the idling point of the isolated two-qubit subsystems L(R) $\Theta_{\rm Id,R(L)}\equiv 0(0)$ (see Appendix D). Similarly to the DTC architecture, the corresponding qubit frequencies of the three-qubit and two-qubit systems are in good agreement with each other. Moreover, we find that $\zeta_{12}$ and $\zeta_{23}$, are almost unchanged from the ones of the isolated two-qubit systems and $\zeta_{13}$ and $\zeta_{ZZZ}$ are negligibly small (about 10 and 0 kHz, respectively). These results show that STCs can suppress the residual couplings even in three-qubit systems. Therefore, we can take the idling point of the three-qubit system to $(\Theta_{\rm ex,L},\Theta_{\rm ex,R})=(\Theta_{\rm Id,L},\Theta_{\rm Id,R})$. However, as in the two-qubit system, their magnitudes shown in Table 5 are about an order of magnitude larger than the corresponding ones of the DTC architecture. These results show that the DTC is superior to the STC in terms of the residual couplings in three-qubit systems, as well as in two-qubit systems.
@@ IV.4.2 Gate performance
Table 6: Average gate fidelities implemented by optimized pulses in the STC architecture.
$\bar{F}_{CZ,12}$ $\bar{F}_{CZ,23}$ $\bar{F}_{\frac{\pi}{2},1}$ $\bar{F}_{\frac{\pi}{2},2}$ $\bar{F}_{\frac{\pi}{2},3}$
Three-qubit system 0.9981 0.9938 0.9980 1.0000 0.9979
Two-qubit system 1.0000 1.0000 ——– ——– ——–
Gate simulation results are summarized in Table  6. We find that 30-ns CZ gates with average fidelities of 0.9999 cannot be achieved in the three-qubit system. This is in contrast to the corresponding fidelities of the two-qubit subsystems, which exceeds 0.9999. The fidelity degradation of the CZ gate due to the scale-up from two qubits to three is about 0.001–0.005, which is roughly 10 times larger than the corresponding one in the DTC architecture. We also find that 10-ns $\pi/2$-pulses with average fidelities over 0.9999 cannot be implemented except for the $\rm{Q_{2}}$. The infidelities of the $\pi/2$ pulses for $\rm{Q_{1}}$ and $\rm{Q_{3}}$, are about 0.002. This is also roughly 10 times larger than the corresponding ones of the system of the DTC.
@@ V Discussion
The above difference in $g_{13}$ arises from the structure of the coupler and the coupling cancellation mechanism. Since the STCs utilizes the relatively large direct capacitances between the NN qubits, $C_{12}$ and $C_{13}$, for canceling residual couplingsyan2018tunable , NN qubits tends to be close to each other. This leads to making NNN qubits close as well, and thus the capacitance between the NNN qubits $C_{13}$ tends to be large. These relatively large qubit-qubit direct capacitances, especially the large $C_{13}$, make $g_{13}$ larger. On the other hand, the DTC does not require $C_{12}$ and $C_{23}$, allowing for greater distances between NN qubits and thus between NNN qubits. This means that in the DTC architecture, it is possible to ideally reduce the direct capacitance between qubits ($C_{12},C_{23},$ and $C_{13}$) and consequently the portion of $g_{13}$ resulting from them to zero.
Note that even if the above direct capacitances are zero, there is still an effective capacitance through the coupler transmons, and hence total $g_{13}$ is still not zero. This effect becomes smaller as the total shunt capacitance of the coupler becomes larger. Since the DTC and the STC are composed of two and one transmons, respectively, a total shunt capacitance in the DTC is typically about twice as large as that in the STC ETH ; STCChinese ; goto2022DTC ; KG2023 . In fact, in our simulations, the total shunt capacitance of each DTC, $C_{44}+C_{55}=C_{66}+C_{77}=160\ {\rm fF}$, is roughly twice as large as the one of each STC, $C_{44{\rm A}}+C_{44{\rm B}}=C_{55{\rm A}}+C_{55{\rm B}}=60\ {\rm fF}$. From this, the DTCs can reduce $g_{13}$ more easily than the STCs.

### v2

@@ III.2.1 Two-qubit gate: CZ gate
To implement $\hat{U}_{CZ,12(23)}$ in the three-qubit system, we fix $\Theta_{\rm ex,R(L)}$ to $\Theta_{\rm Id,R(L)}$ and apply a optimized flux pulse $\Theta_{\rm ex,L(R)}(t)$ (see Appendix G.1). We can successfully implement 30-ns CZ gates with average fidelities over 99.99% for both the qubit pairs, as summarized in Table  3. Comparing these fidelities with the corresponding ones in the isolated two-qubit subsystems, we find that the degradation of the CZ gate fidelity due to the increasing number of qubits from two to three is negligible in the 4-digit precision.
@@ III.2.2 Single-qubit gate: $\pi/2$ pulse
Next, we consider the single-qubit gate as well. Our targets are individual and simultaneous implementations of $10$-ns $\pi/2$ pulses. It leads to the following $\hat{U}_{\rm id}$:
$\hat{U}_{\pi/2,\{1\}}$ $=\hat{U}_{\pi/2}\otimes\hat{I}\otimes\hat{I},$ (14)
$\hat{U}_{\pi/2,\{2\}}$ $=\hat{I}\otimes\hat{U}_{\pi/2}\otimes\hat{I},$ (15)
$\hat{U}_{\pi/2,\{3\}}$ $=\hat{I}\otimes\hat{I}\otimes\hat{U}_{\pi/2},$ (16)
$\hat{U}_{\pi/2,\{1,2\}}$ $=\hat{U}_{\pi/2}\otimes\hat{U}_{\pi/2}\otimes\hat{I},$ (17)
$\hat{U}_{\pi/2,\{2,3\}}$ $=\hat{I}\otimes\hat{U}_{\pi/2}\otimes\hat{U}_{\pi/2},$ (18)
$\hat{U}_{\pi/2,\{1,3\}}$ $=\hat{U}_{\pi/2}\otimes\hat{I}\otimes\hat{U}_{\pi/2},$ (19)
$\hat{U}_{\pi/2,\{1,2,3\}}$ $=\hat{U}_{\pi/2}\otimes\hat{U}_{\pi/2}\otimes\hat{U}_{\pi/2},$ (20)
@@
$\hat{U}_{\pi/2}=\frac{1}{\sqrt{2}}\begin{pmatrix}1&-i\\ -i&1\\ \end{pmatrix}.$ (21)
The average gate fidelity of $\hat{U}_{\pi/2,k}$ is denoted by $\bar{F}_{\pi/2,k}$ for $k\in\{\{1\},\{2\},\{3\},\{1,2\},\{2,3\},\{1,3\},\{1,2,3\}\}$.
To implement $\hat{U}_{\pi/2,k}$, we fix $\Theta_{\rm ex,\mu}$ to $\Theta_{\rm Id,\mu}$ and apply a driving voltage to ${\rm Q}_{i}$ for $i\in k$. This operation is described by the following Hamiltonian (see Appendix A for the derivation):
$\hat{H}_{V}$ $=4\hbar\hat{\mathbf{n}}^{\rm T}W\hat{\mathbf{n}}+\sum_{i\in k}\alpha_{i}\sum_{j=1}^{7}\hbar W_{ij}\hat{n}_{j}+\hat{U}(\Theta_{\rm Id,L},\Theta_{\rm Id,R}),$ (22)
where $\alpha_{i}$ is a dimensionless parameter controlled by $V_{i}$ (see Appendix A). Applying optimized $\alpha_{i}(t)$ (see Appendix G.2), we can implement individual and simultaneous 10-ns $\pi/2$ pulses with average gate fidelities over 99.99% as summarized in Table  3.
@@
In this section, we have shown that DTCs can not only reduce undesired residual couplings to several kHz but also enable implementations of fast CZ gates and individual and simultaneous $\pi/2$ pulses with average fidelities over 99.99%, in the three-qubit system as well as in isolated two-qubit systems. These results indicate that the DTC architecture is scalable for highly detuned fixed-frequency qubits.
@@ IV.2 Hamiltonian of the superconducting circuit model
Note that almost all previous theoretical analyses of the STC reported in the literature have been based on the effective model yan2018tunable ; PRXSTC ; RiggettiSTC ; STCChinese ; IBMSTC ; YanSTCCZ ; TheoriticalSTCcrosstalk2022 . Unlike them, we use the circuit model for higher accuracy, as well as the DTCs. To the best of our knowledge, this is the first estimation of $ZZ$ coupling between NNN qubits and $ZZZ$ coupling and also the first gate simulation based on the circuit model for systems with the STCs.
@@ IV.3 Comparing conditions
The preferable parameter setting for the comparison between the STC and DTC architectures is non-trivial. This is because the STC and the DTC obey different operating principles, resulting in different parameter values desired for high performance. Here, in order to make the comparison as fair as possible, we set the parameters as shown in Table 4 , considering the following points. (1) We set $\omega_{i}$ and $C_{ii}$ for $i\in\{1,2,3\}$ to be the same as in Table 1, which means that qubit frequencies and anharmonicities are almost the same as in the DTC case. (2) Under this condition, we set $C_{ij}$ that exists in the two-qubit subsystems and $I_{c4{\rm B}(c5{\rm B})}/I_{c4\rm{A}(c5{\rm A})}$ to the almost same values as the corresponding ones in Ref. ETH, , and we also set $\omega_{4(5)}$ to the almost same value as the coupler idle frequency in Ref. ETH, . Reference ETH, experimentally implemented a high-performance CZ gate (with a gate time of 38 ns and a fidelity of 97.9%), using a parameter setting close to the above condition (1) , where qubits frequencies are 5.038 and 5.400 GHz and both qubits capacitances are 77.8 fF. Moreover, they also showed that the magnitude of the residual $ZZ$ coupling can be suppressed to 60–80 kHz by an STC, which is known to be one of the most $ZZ$-coupling suppressed results for highly detuned qubits where a fast CZ gate can be implemented. We will show that our numerical simulation reproduces their results well. (3) As for the rest $C_{ij}$ that does not exist in the two-qubit subsystems (e.g. $C_{13},C_{15}$, and $C_{45}$), we choose feasible values estimated by electromagnetic simulator ANSYS Q3DAnsys assuming the scaled-up design of the circuit in Ref. ETH, (see Fig. 5).
Figure 5: The scaled-up design of the circuit in Ref. ETH, . Its capacitances estimated by electromagnetic simulator ANSYS Q3DAnsys are summarized in Table 4.
@@ IV.4 Numerical results
Figure 6: Energy levels and $ZZ$ coupling strengths of the systems in Fig. 4. (a) and (b) show eigenfrequencies $\omega_{Q_{1},Q_{2},Q_{3},C_{4},C_{5}}$ as functions of $\Theta_{\rm ex,L}$ and $\Theta_{\rm ex,R}$, respectively. The qubit frequencies of the three-qubit system, $\omega_{Q_{1},Q_{2},Q_{3}}$ $(Q_{i}\in\{0,1\})$, are highlighted by being colored and bold, and the ones of the two-qubit system, $\omega_{Q_{i},Q_{j}}^{\mu}$ ($Q_{i},Q_{j}\in\{0,1\}$ and $\mu\in\{{\rm L,R}\})$ (see Appendix D), are shown by colored scatter plots. Eigenfrequencies $\omega_{Q_{1},Q_{2},Q_{3},C_{4},C_{5}}$ other than the above are represented by black thin curves. (c) and (d) represent $ZZ$ and $ZZZ$ couplings as functions of $\Theta_{\rm ex,L}$ and $\Theta_{\rm ex,R}$, respectively. Solid curves represent the ones of the three-qubit system, and scatter plots represent $\zeta_{12(23)}$ in the system L(R). The Orange (blue) dotted horizontal line is $\zeta_{\rm Id,23(12)}$ in the system R(L). Note that $\Theta_{\rm ex,R}$ in the left column and $\Theta_{\rm ex,L}$ in the right column are fixed to $\Theta_{\rm Id,R}$ and $\Theta_{\rm Id,L}$, respectively.
@@
IV.4.1 Idle point, $ZZ$ couplings, and $ZZZ$ coupling
Figure 6(a)[6(b)] shows $\omega_{Q_{1},Q_{2},Q_{3},C_{4},C_{5}}$ as functions of $\Theta_{\rm L(R)}$ when $\Theta_{\rm ex,R(L)}$ is fixed to the idle point of the isolated two-qubit subsystems L(R) $\Theta_{\rm Id,R(L)}\equiv 0(0)$ (see Appendix D). The qubit frequencies of the three-qubit system, $\omega_{Q_{1},Q_{2},Q_{3}}$ $(Q_{i}\in\{0,1\})$, are highlighted by being colored and bold, and the ones of the two-qubit system, $\omega_{Q_{i},Q_{j}}^{\mu}$ ($Q_{i},Q_{j}\in\{0,1\}$ and $\mu\in\{{\rm L,R}\})$ are shown by scatter plots. All of them are needed for calculations of $ZZ$ and $ZZZ$ couplings below. Similarly to the DTC architecture, we find that the corresponding qubit frequencies of the three-qubit and two-qubit systems are in good agreement with each other. The black thin curves represent the states other than the computational ones. They are necessary for accounting leakage errors during gate simulations.
Figure 6(c)[6(d)] shows the $ZZ$ and $ZZZ$ couplings. Solid curves represent the ones of the three-qubit system, and scatter plots represent the ones of the two-qubit system. We find that $\zeta_{12}$ and $\zeta_{23}$, are almost unchanged from the ones of the isolated two-qubit systems and $\zeta_{13}$ and $\zeta_{ZZZ}$ are negligibly small (about 10 and 0 kHz, respectively). These results show that STCs can suppress the residual couplings even in three-qubit systems. Therefore, we can take the idle point of the three-qubit system to $(\Theta_{\rm ex,L},\Theta_{\rm ex,R})=(\Theta_{\rm Id,L},\Theta_{\rm Id,R})$. However, as in the two-qubit system, their magnitudes shown in Table 5 are about an order of magnitude larger than the corresponding ones of the DTC architecture. These results show that the DTC is superior to the STC in terms of the residual couplings in three-qubit systems, as well as in two-qubit systems.
Table 6: Average gate fidelities of 30-ns CZ gates and 10-ns individual and simultaneous $\pi/2$ pulses. They are implemented by optimized pulses in the STC architecture.
Average gate fidelity (%) $\bar{F}_{CZ,12}$ $\bar{F}_{CZ,23}$ $\bar{F}_{\frac{\pi}{2},\{1\}}$ $\bar{F}_{\frac{\pi}{2},\{2\}}$ $\bar{F}_{\frac{\pi}{2},\{3\}}$ $\bar{F}_{\frac{\pi}{2},\{1,2\}}$ $\bar{F}_{\frac{\pi}{2},\{2,3\}}$ $\bar{F}_{\frac{\pi}{2},\{1,3\}}$ $\bar{F}_{\frac{\pi}{2},\{1,2,3\}}$
Three-qubit system 99.81 99.38 99.80 100.00 99.79 99.80 99.79 99.63 99.62
Two-qubit system 100.00 100.00 ——– ——– ——– ——– ——– ——– ——–
@@ IV.4.2 Gate performance
Gate simulation results are summarized in Table  6. We find that 30-ns CZ gates with average fidelities of 99.99% cannot be achieved in the three-qubit system. This is in contrast to the corresponding fidelities of the two-qubit subsystems, which exceeds 99.99%. The fidelity degradation of the CZ gate due to the scale-up from two qubits to three is about 0.2%–0.6%, which is roughly 10 times larger than the corresponding one in the DTC architecture. We also find that 10-ns $\pi/2$-pulses with average fidelities over 99.99% cannot be implemented except for the $\rm{Q_{2}}$. The infidelities of the $\pi/2$ pulses for $\rm{Q_{1}}$ and $\rm{Q_{3}}$, are about 0.2%. This is also roughly 10 times larger than the corresponding ones of the system of the DTC.
@@ V Discussion
The above difference in $g_{13}/(2\pi)$ (0.13 MHz and 1.62 MHz for the DTC and the STC architecture, respectively) arises from the structure of the coupler and the coupling cancellation mechanism. Since the STCs utilize the relatively large direct capacitances between the NN qubits, $C_{12}$ and $C_{23}$, for canceling residual couplingsyan2018tunable , NN qubits tend to be close to each other. This makes NNN qubits close as well, and thus the capacitance between the NNN qubits $C_{13}$ tends to be large (0.03 fF). On the other hand, the DTC does not require $C_{12}$ and $C_{23}$, allowing for greater distances between NN qubits and hence between NNN qubits, leading to a smaller $C_{13}$ (0.003 fF). This difference in $C_{13}$ is a major reason for the difference in $g_{13}$.
Note that even if $C_{13}$ is zero, there is still an effective capacitance, and hence the $g_{13}$ is not zero: 0.04 MHz and 0.71 MHz for the DTC and the STC architecture, respectively. The one in the STC still remains about a half of the original value.
In an ideal limit where all parasitic capacitances are zero, $g_{13}/(2\pi)$ of the DTC and STC become 0.00 MHz and 0.38 MHz, respectively. This means that in the DTC architecture, $g_{13}$ arises through parasitic capacitances and vanishes in the ideal situation. On the other hand, even in this ideal situation, $g_{13}$ in the STC architecture is about a half of that before taking the limit.
The remaining $g_{13}$ in the STC architecture is determined by two factors. One of them is the direct couplings between NN qubits, namely $C_{12}$ and $C_{23}$. In the STC architecture, they are large for canceling residual couplings as mentioned above, and thus this contribution does not vanish even in the above ideal situation with no parasitic capacitances. On the other hand, the DTC does not rely on $C_{12}$ and $C_{13}$, that is, $C_{12}$ and $C_{23}$ are parasitic capacitances for the DTC architecture, and hence the contribution is negligible.
The other factor is the indirect couplings through a coupler transmon. This contribution can be estimated by considering the case where $C_{12}$ and $C_{23}$ are zero, resulting in 0.11 MHz for the STC architecture. In the DTC architecture, corresponding contribution is negligible because the coupling between the two transmons inside DTC L(R) is nearly disconnected due to the small parasitic capacitance $C_{45}$($C_{67}$).
In summary, the higher performance of the DTC architecture than the STC one mainly comes from the fact that the DTC has the advantage that it can reduce the NNN coupling, $g_{13}$, compared to the STC from various points.